Calculate the pH at the equivalence volume for the titration of 25.0 mL of 0.350 M acetic acid with 0.350 M sodium hydroxide. The K a for acetic acid is 1.8 * 10 ^ - 5 .
Answer #1

pKa = - log Ka = - log (1.8 * 10^-5) = 4.745

volume of NaOH solution needed at equivalence point = 25.0 mL * 0.350 M / 0.350 M = 25.0 ml

and

total volume of solution = (25.0 + 25.0) = 50.0 ml

at equivalence point,

CH3COOH + NaOH ..........> CH3COONa + H2O

[salt] = 25.0 ml * 0.350 M / 50.0 ml = 0.175 M

so

salt hydrolysis occurs.

pH = 1/2 * [pKw + pKa + log C]

or

pH = 1/2 * [14 + 4.745 + log (0.175)]

or

pH = 8.994

or

pH = 9.0 (answer)